OpenStudy (anonymous):

A car that sold for $11,800 has a sale tax of $767. How much does a car cost if its sales tax is $637

OpenStudy (texaschic101):

11,800/767 = x/637 cross multiply 767x = (11800)(637) 767x = 7516600 x = 7516600/767 x = 9800 check.. 11,800/767 = 9800/637 cross multiply (767)(9800) = (11,800)(637) 7516600 = 7516600 (correct) The car would cost 9800 to have a sales tax of 637